A ball rolls off a 108 cm high table at 5.6 m/s. What is its speed when it hits the ground, in m/s?

A baseball is hit at an angle of 26.3° and velocity of 26.2 m/s on level ground. A stiff head wind causes a horizontal acceleration of 1.97 m/s2 opposing the ball's motion. How far in meters does the ball travel on level ground?

You are the quarterback ready to pass the ball to a receiver already 23.5 m down the field (from you). He is running straight away at a constant 6.31 m/s, and you will release the ball at an angle of 45.9°. With what speed (in m/s) should you throw the ball?
1
Expert's answer
2019-09-06T09:44:10-0400

1.

"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(1.08)}{9.8}}=0.469\ s"

"v=\sqrt{(v_x)^2+(v_y)^2}=\sqrt{(v_x)^2+(gt)^2}"

"v=\sqrt{(5.6)^2+((9.8)(0.469))^2}=7.2\frac{m}{s}"

2.


"t=\frac{2v\sin{26.3°}}{g}=\frac{2(26.2)\sin{26.3°}}{9.8}=2.369\ s"

"s=v\cos{26.3°}t-0.5at^2"

"s=26.2\cos{26.3°}(2.369)-0.5(1.97)(2.369)^2=50.1\ m"

3.


"x=vt+s"

"t=\frac{2V\sin{45.9°}}{g}"

"x=V \cos{45.9°}t"

"V \cos{45.9°}t=vt+s"

"V \cos{45.9°}\frac{2V\sin{45.9°}}{g}=v\frac{2V\sin{45.9°}}{g}+s"

"V \cos{45.9°}\frac{2V\sin{45.9°}}{9.8}=6.31\frac{2V\sin{45.9°}}{9.8}+23.5"

"V=20.4\frac{m}{s}"